If Leah is 6 years older than sue, and John is 5 years older than Leah, and the total of their ages is 41. Then how old is sue?
I don't really need the answer I need help with the formula to get the answer, can some one show how to do this?

Answer by ankor@dixie-net.com(22740)

If Leah is 6 years older than sue, and John is 5 years older than Leah, and the total of their ages is 41.
Then how old is sue?
:
Let L = Leah's age
let S = Sue's age
Let J = John's age
:
WRite an equation for each statement, arrange to get S and J in terms of L:
:
"If Leah is 6 years older than sue,"
L = S + 6
S = L - 6
:
" John is 5 years older than Leah,"
J = L + 5
:
" the total of their ages is 41."
L + S + J = 41
:
Replace S with (L-6), replace J with (L+5), solve for L, then find S & J
L + (L-6) + (L+5) = 41
:
Think you can handle it now?